# Chapter contributed to the volume “Time and timelessness in fundamental physics and cosmology”, edited by S. De Bianchi, M. Forgione, and L. Marongiu, Fundamental Theories of Physics series (Springer)]

### v1

@@ Abstract
Different approaches to quantum gravity, such as loop quantum cosmology and group field theory, predict the resolution of the initial cosmological singularity via a ‘bounce’: a regular spacetime region that connects the expanding branch of the universe to a contracting branch. The cosmological arrow of time, which by definition points in the direction of cosmic expansion, is reversed at the bounce. Nonetheless, it is still possible to discriminate between the two branches by considering different arrows, as defined for instance by the growth of perturbations. After reviewing general aspects of the time arrow problem in cosmology, we examine the properties of different arrows of time in bouncing cosmologies, focusing on the loop quantum cosmology bounce as a case study. We also present a new exact solution to the effective Friedmann equations of loop quantum cosmology with pressureless dust and a cosmological constant, which is a simplified version of the $\Lambda$CDM bounce scenario, where these issues can be examined in detail.
@@ I Arrows of time and their cosmological origin
Further arrows of time are also defined by, e.g., state vector reduction in the measurement of a quantum system, diffusive and radiative processes in astrophysics, biological processes, and so on Ellis (2013).
Although conceptually distinct, the different arrows of time are not independent of one another. For instance, the psychological arrow clearly depends on the thermodynamical one, as biological systems obey the laws of thermodynamics. The electrodynamical arrow is determined by a boundary condition of no incoming radiation in the past, which in the Wheeler-Feynman absorber theory is traced back to the thermodynamic properties of the absorbing medium Hawking (1985); Zeh (2011); Forgione (2022). More in general, if the physical descriptions of systems are organized hierarchically, ranging from the smallest to the largest scales through successive levels of emergence, it is possible to identify relations between different levels that proceed in both a bottom-up and a top-down manner Ellis (2013).
@@
Alternatively, attempts have been made to explain the low entropy of the observable universe starting from a random initial state as the result of a dynamical process, such as inflation. However, special initial conditions must be assumed at the onset of inflation in order to produce an inflating patch that evolves to a universe like ours; thus, the issue of the exceptionality of the initial state is just shifted to the pre-inflationary epoch Wald (2006); Earman (2006); Davies (2014). Yet another approach to explain the initial conditions of the universe is to appeal to quantum cosmology Vilenkin (1983); Hartle and Hawking (1983); Feldbrugge et al. (2017), where the problem translates into the identification of appropriate boundary conditions in superspace, or, equivalently, of specific integration contours in the path integral over geometries. There are also proposals where entropy grows symmetrically both to the past and to the future of a given low-entropy spatial hypersurface Aguirre and Gratton (2003); Carroll and Chen (2004); Barbour et al. (2014); Vilenkin (2013). Two-times high-entropy boundary conditions can also lead to intermediate low entropy states, as shown in Ref. Deutsch and Aguirre (2022).
@@ II The LQC bounce
It is well-known that classical general relativity predicts its own breakdown at spacetime singularities Hawking and Penrose (1970). There is a widespread expectation that such singularities must be resolved in a theory of quantum gravity. In the cosmological case, this idea has been realized concretely in the framework of non-perturbative approaches to quantum gravity, such as loop quantum cosmology (LQC) Ashtekar and Singh (2011) and group field theory (GFT) Oriti (2006); Oriti et al. (2016), where the initial cosmological “Big Bang” singularity is replaced by a “Big Bounce” (see also Oriti’s contribution in this volume Oriti (2024) and the reviews Pithis and Sakellariadou (2019); Gabbanelli and De Bianchi (2021)). In these models, a nonsingular bounce is obtained as a result of quantum gravity effects, which cause the gravitational interaction to turn into a repulsive force at high energy densities, typically of the order of the Planck density. The bounce represents a smooth spacetime region that connects the expanding branch of the universe to a past contracting branch. A similar scenario has been realized also in Pre-Big Bang cosmology in the framework of string theory Gasperini and Veneziano (2003, 2023); Veneziano (1999). Besides quantum gravity, singularity resolution has also been studied in the framework of modified gravity, where early proposals were based on the implementation of the so-called limiting curvature hypothesis Markov (1982); Mukhanov and Brandenberger (1992), and bouncing cosmologies have been concretely realized more recently in different classes of theories (including, e.g. scalar-tensor theories, massive gravity, and higher-order modifications of general relativity, see Easson et al. (2011); Cai et al. (2012a, b); Ijjas and Steinhardt (2016); Chamseddine and Mukhanov (2017); de Cesare (2019a); de Cesare and Wilson-Ewing (2019); Ilyas et al. (2020); de Cesare (2019b); Ilyas et al. (2020) and references therein). Bouncing models resolve the classical horizon problem of standard cosmology, and may represent either an alternative to the inflationary mechanism for the generation of primordial cosmological fluctuations or a complement to inflation that includes the effects of Planck scale physics Brandenberger and Peter (2017); Agullo et al. (2023).
In this section we focus specifically on the LQC bounce. After reviewing the effective cosmological dynamics, we will consider a concrete model where the energy content consists of pressureless dust and a cosmological constant, and derive a novel exact solution of the LQC effective Friedmann equations. We will show that no physically meaningful time asymmetry exists at the background level, i.e. both the cosmological and the thermodynamical arrows of time are collinear and point away from the bounce. The dynamics of perturbations and the resulting arrow of time will be studied in the following section.
@@
For the sake of concreteness, we consider a model with a cosmological constant $\Lambda>0$ and a pressureless fluid $\rho=\rho_{\rm m}+\Lambda/(8\pi G)$, with $\rho_{\rm m}=\rho_{o}(a_{o}/a)^{3}$, where $a_{o}$ is the scale factor at the bounce. The presence of a positive cosmological constant ensures that the universe is asymptotically de Sitter, and therefore one can define cosmological event horizons (as well as their entropy). The present model is a simplified version of the $\Lambda$CDM bounce scenario in LQC, which includes also the contribution of radiation (which dominates over dust in the early universe) and has been studied earlier in Refs. Cai and Wilson-Ewing (2015); Cai et al. (2016). In turn, the $\Lambda$CDM bounce is a generalization of the LQC matter bounce scenario Wilson-Ewing (2013) (for a recent review see Ref. Agullo et al. (2023)). Using Eq. (2.1a) at the bounce, we obtain $\rho_{o}+\Lambda/(8\pi G)=\rho_{c}$, which can be used to eliminate $\rho_{o}$. The model is exactly solvable, with scale factor given by [^7] A generalization of this solution can be obtained in a model with a cosmological constant and a fluid with arbitrary equation of state $w$ (in which case the energy density is $\rho=\rho_{w}+\Lambda/(8\pi G)$, with $\rho_{w}=\rho_{o}(a_{o}/a)^{3(w+1)}$ ). Also in this case there exists an exact solution of the LQC effective Friedmann equation (2.1a) with a very similar structure as (2.3), given by $a(t)=a_{o}\left[1+\frac{4\pi G\rho_{c}}{\Lambda}\bigg{(}\cosh\Big{(}3(1+w)H_{\infty}t\Big{)}-1\bigg{)}\right]^{\frac{1}{3(1+w)}}$.
@@
where we defined $H_{\infty}\equiv\sqrt{\Lambda\left(1-\Lambda/(8\pi G\rho_{c})\right)/3~{}}$ To the best of our knowledge, this is a new exact solution (2.3) of the effective Friedmann equations in LQC.
@@ III Time arrow of perturbations in the $\Lambda$CDM bounce scenario
We consider the $\Lambda$CDM bounce scenario of Refs. Cai and Wilson-Ewing (2015); Cai et al. (2016) and review the evolution of scalar perturbations in their model. As in the simplified model considered in the previous section, also the $\Lambda$CDM bounce is symmetric. The difference between the two models lies in the fact is that the $\Lambda$CDM bounce also takes into account the contribution of radiation, which dominates the energy density of the universe at the bounce. Moreover, there is an intermediate era of radiation domination between the bounce and the matter dominated era. At late times, as well as in the distant past, radiation becomes sub-leading compared to non-relativistic matter and dark energy, and therefore the evolution of the scale factor in this regime is well approximated by (2.3). In both models, general relativity is recovered in the large $|t|$ limit.
@@
The evolution of perturbations in this model has been studied using both analytical and numerical methods in Ref. Cai and Wilson-Ewing (2015) (to which we refer the reader for more details). After the bounce and at horizon re-entry, the leading contribution to the curvature perturbation is a constant mode. The scalar power spectrum reads
@@ IV Discussion
We have seen that a number of different arrows of time can be introduced in bouncing cosmologies. At the background level, the cosmological arrow of time always points in the direction of expansion, and therefore away from the bounce. It is also possible to define a thermodynamical arrow of time, pointing in the direction of horizon entropy growth. However, if the bounce is symmetric also the background thermodynamical arrow is unable to introduce any distinction between the two branches. We examined this issue in detail for a newly derived exact solution of the effective Friedmann equations of LQC, describing a bouncing cosmology with pressureless dust and a positive cosmological constant. This simple example shows that, in general, the analysis of the evolution of the cosmological background is not sufficient to draw any reliable conclusions about the arrow of time. We have then examined the evolution of scalar perturbations and discussed how the time arrow of inhomogeneities growth originates from initial conditions that break time-reversal symmetry. A specific choice of (vacuum) initial conditions is made in the distant past, which amounts to a Past Hypothesis. This plays an analogous role to Penrose’s Weyl hypothesis in standard cosmology in selecting a special initial state for the universe.
@@
Lastly, if spacetime emerges dynamically from a pre-geometric structure, as argued for instance within the context of GFT and in the causal sets approach to quantum gravity (see contributions by Oriti and Forgione in this volume, and Ref. De Bianchi and Gabbanelli (2023)), the problem of defining an ‘initial state’ for the universe may require a reframing that makes no reference to the emergent spacetime geometry. Within this framework, it would be interesting to investigate whether the initial conditions of the universe could be explained from first principles, for instance by means of a dynamical mechanism, or if a Past Hypothesis (i.e., a law-like prescription similar to the Weyl curvature hypothesis) is still required.
@@ Acknowledgments
I would like to thank Edward Wilson-Ewing for helpful comments and Silvia De Bianchi for the invitation to contribute to the volume “Time and timelessness in fundamental physics and cosmology”. The author acknowledges support from INFN, iniziative specifiche QUAGRAP and GeoSymQFT.

### v2

@@ Abstract
Different approaches to quantum gravity, such as loop quantum cosmology and group field theory, predict the resolution of the initial cosmological singularity via a ‘bounce’: a regular spacetime region that connects the expanding branch of the universe to a contracting branch. The cosmological arrow of time, which by definition points in the direction of cosmic expansion, is reversed at the bounce. Nonetheless, it is still possible to discriminate between the two branches by considering different arrows, as defined for instance by the growth of perturbations. After reviewing general aspects of the time arrow problem in cosmology, we examine the properties of different arrows of time in bouncing cosmologies, focusing on the loop quantum cosmology bounce as a case study. These issues are examined in detail for an exact solution to the effective Friedmann equations of loop quantum cosmology with pressureless dust and a cosmological constant, which is a simplified version of the $\Lambda$CDM bounce scenario.
@@ I Arrows of time and their cosmological origin
Further arrows of time are also defined by, e.g., state vector reduction in the measurement of a quantum system, diffusive and radiative processes in astrophysics, biological processes, and so on Ellis (2013). Although conceptually distinct, the different arrows of time are not independent of one another. For instance, the psychological arrow clearly depends on the thermodynamical one, as biological systems obey the laws of thermodynamics. The electrodynamical arrow is determined by a boundary condition of no incoming radiation in the past, which in the Wheeler-Feynman absorber theory is traced back to the thermodynamic properties of the absorbing medium Hawking (1985); Zeh (2011); Forgione (2022). More in general, if the physical descriptions of systems are organized hierarchically, ranging from the smallest to the largest scales through successive levels of emergence, it is possible to identify relations between different levels that proceed in both a bottom-up and a top-down manner Ellis (2013).
@@
Alternatively, attempts have been made to explain the low entropy of the observable universe starting from a random initial state as the result of a dynamical process, such as inflation. However, special initial conditions must be assumed at the onset of inflation in order to produce an inflating patch that evolves to a universe like ours; thus, the issue of the exceptionality of the initial state is just shifted to the pre-inflationary epoch Wald (2006); Earman (2006); Davies (2014). Yet another approach to explain the initial conditions of the universe is to appeal to quantum cosmology Vilenkin (1983); Hartle and Hawking (1983); Feldbrugge et al. (2017), where the problem translates into the identification of appropriate boundary conditions in superspace, or, equivalently, of specific integration contours in the path integral over geometries. There are also proposals where entropy grows symmetrically both to the past and to the future of a given low-entropy spatial hypersurface Aguirre and Gratton (2003); Carroll and Chen (2004); Vilenkin (2013). [^6] An alternative proposal with bidirectional arrows of time representing growth of complexity was made in Barbour et al. (2014). Two-times high-entropy boundary conditions can also lead to intermediate low entropy states, as shown in Ref. Deutsch and Aguirre (2022).
@@ II The LQC bounce
It is well-known that classical general relativity predicts its own breakdown at spacetime singularities Hawking and Penrose (1970). There is a widespread expectation that such singularities must be resolved in a theory of quantum gravity. In the cosmological case, this idea has been realized concretely in the framework of non-perturbative approaches to quantum gravity, such as loop quantum cosmology (LQC) Ashtekar and Singh (2011) and group field theory (GFT) Oriti (2006); Oriti et al. (2016), where the initial cosmological “Big Bang” singularity is replaced by a “Big Bounce” (see also Oriti’s contribution in this volume Oriti (2024) and the reviews Pithis and Sakellariadou (2019); Gabbanelli and De Bianchi (2021)). [^7] See also Ref. Barca et al. (2021). In these models, a nonsingular bounce is obtained as a result of quantum gravity effects, which cause the gravitational interaction to turn into a repulsive force at high energy densities, typically of the order of the Planck density. The bounce represents a smooth spacetime region that connects the expanding branch of the universe to a past contracting branch. A similar scenario has been realized also in Pre-Big Bang cosmology in the framework of string theory Gasperini and Veneziano (2003, 2023); Veneziano (1999), see also Ref. Conzinu et al. (2023). Besides quantum gravity, singularity resolution has also been studied in the framework of modified gravity, where early proposals were based on the implementation of the so-called limiting curvature hypothesis Markov (1982); Mukhanov and Brandenberger (1992), and bouncing cosmologies have been concretely realized more recently in different classes of theories (including, e.g. scalar-tensor theories, massive gravity, and higher-order modifications of general relativity, see Easson et al. (2011); Cai et al. (2012a, b); Ijjas and Steinhardt (2016); Chamseddine and Mukhanov (2017); de Cesare (2019a); de Cesare and Wilson-Ewing (2019); Ilyas et al. (2020); de Cesare (2019b); Ilyas et al. (2020) and references therein). Bouncing models resolve the classical horizon problem of standard cosmology, and may represent either an alternative to the inflationary mechanism for the generation of primordial cosmological fluctuations or a complement to inflation that includes the effects of Planck scale physics Brandenberger and Peter (2017); Agullo et al. (2023).
In this section we focus specifically on the LQC bounce. After reviewing the effective cosmological dynamics, we will consider a concrete model where the energy content consists of pressureless dust and a cosmological constant, and derive an exact solution of the LQC effective Friedmann equations. We will show that no physically meaningful time asymmetry exists at the background level, i.e., both the cosmological and the thermodynamical arrows of time are collinear and point away from the bounce. The dynamics of perturbations and the resulting arrow of time will be studied in the following section.
@@
For the sake of concreteness, we consider a model with a cosmological constant $\Lambda>0$ and a pressureless fluid $\rho=\rho_{\rm m}+\Lambda/(8\pi G)$, with $\rho_{\rm m}=\rho_{o}(a_{o}/a)^{3}$, where $a_{o}$ is the scale factor at the bounce. The presence of a positive cosmological constant ensures that the universe is asymptotically de Sitter, and therefore one can define cosmological event horizons (as well as their entropy). The present model is a simplified version of the $\Lambda$CDM bounce scenario in LQC, which includes also the contribution of radiation (which dominates over dust in the early universe) and has been studied earlier in Refs. Cai and Wilson-Ewing (2015); Cai et al. (2016). In turn, the $\Lambda$CDM bounce is a generalization of the LQC matter bounce scenario Wilson-Ewing (2013) (for a recent review see Ref. Agullo et al. (2023)). Using Eq. (2.1a) at the bounce, we obtain $\rho_{o}+\Lambda/(8\pi G)=\rho_{c}$, which can be used to eliminate $\rho_{o}$. The model is exactly solvable, with scale factor given by [^9] A generalization of this solution can be obtained in a model with a cosmological constant and a fluid with arbitrary equation of state $w$ (in which case the energy density is $\rho=\rho_{w}+\Lambda/(8\pi G)$, with $\rho_{w}=\rho_{o}(a_{o}/a)^{3(w+1)}$ ). Also in this case there exists an exact solution of the LQC effective Friedmann equation (2.1a) with a very similar structure as (2.3), given by $a(t)=a_{o}\left[1+\frac{4\pi G\rho_{c}}{\Lambda}\bigg{(}\cosh\Big{(}3(1+w)H_{\infty}t\Big{)}-1\bigg{)}\right]^{\frac{1}{3(1+w)}}$, see Ref. Chinaglia et al. (2017).
@@
where we defined $H_{\infty}\equiv\sqrt{\Lambda\left(1-\Lambda/(8\pi G\rho_{c})\right)/3~{}}$ This solution of the effective Friedmann equations in LQC has been obtained earlier in Ref. Chinaglia et al. (2017).
@@ III Time arrow of perturbations in the $\Lambda$CDM bounce scenario
We consider the $\Lambda$CDM bounce scenario of Refs. Cai and Wilson-Ewing (2015); Cai et al. (2016) and review the evolution of scalar perturbations in their model. As in the simplified model considered in the previous section, also the $\Lambda$CDM bounce is symmetric. The difference between the two models lies in the fact that the $\Lambda$CDM bounce also takes into account the contribution of radiation, which dominates the energy density of the universe at the bounce. Moreover, there is an intermediate era of radiation domination between the bounce and the matter dominated era. At late times, as well as in the distant past, radiation becomes sub-leading compared to non-relativistic matter and dark energy, and therefore the evolution of the scale factor in this regime is well approximated by (2.3). In both models, general relativity is recovered in the large $|t|$ limit.
@@
The evolution of perturbations in this model has been studied using both analytical and numerical methods in Ref. Cai and Wilson-Ewing (2015), to which we refer the reader for more details. After the bounce and at horizon re-entry, the leading contribution to the curvature perturbation is a constant mode. The scalar power spectrum reads
@@ IV Discussion
We have seen that a number of different arrows of time can be introduced in bouncing cosmologies. At the background level, the cosmological arrow of time always points in the direction of expansion, and therefore away from the bounce. It is also possible to define a thermodynamical arrow of time, pointing in the direction of horizon entropy growth. However, if the bounce is symmetric also the background thermodynamical arrow is unable to introduce any distinction between the two branches. We examined this issue in detail for an exact solution of the effective Friedmann equations of LQC, describing a bouncing cosmology with pressureless dust and a positive cosmological constant. This simple example shows that, in general, the analysis of the evolution of the cosmological background is not sufficient to draw any reliable conclusions about the arrow of time. We have then examined the evolution of scalar perturbations and discussed how the time arrow of inhomogeneities growth originates from initial conditions that break time-reversal symmetry. A specific choice of (vacuum) initial conditions is made in the distant past, which amounts to a Past Hypothesis. This plays an analogous role to Penrose’s Weyl hypothesis in standard cosmology in selecting a special initial state for the universe.
@@
Lastly, if spacetime emerges dynamically from a pre-geometric structure, as argued for instance within the context of GFT and in the causal sets approach to quantum gravity (see contributions by Oriti Oriti (2024) and Forgione Forgione (2024) in this volume, and Ref. De Bianchi and Gabbanelli (2023)), the problem of defining an ‘initial state’ for the universe may require a reframing that makes no reference to the emergent spacetime geometry. Within this framework, it would be interesting to investigate whether the initial conditions of the universe could be explained from first principles, for instance by means of a dynamical mechanism, or if a Past Hypothesis (i.e., a law-like prescription similar to the Weyl curvature hypothesis) is still required.
@@ Acknowledgments
I am grateful to Silvia De Bianchi for the invitation to contribute to the volume “Time and Timelessness in Fundamental Physics and Cosmology”, and would like to thank Edward Wilson-Ewing and Sergio Zerbini for helpful comments. The author acknowledges support from INFN, iniziative specifiche QUAGRAP and GeoSymQFT.